### v1

# Quantum criticality and universality in stationary state of long-range Kitaev model

Abstract
We investigate the signature of quantum criticality in the long-time stationary state of the long-range Kitaev chain by performing various quench protocols. In this model, the pairing interaction decays with distance according to a power law with exponent $\alpha$. Using quantum information-theoretic measures, such as mutual information and logarithmic negativity, we show that, irrespective of the values of $\alpha$, critical-to-critical quench displays quantum criticality even in the stationary state. Remarkably, in the presence of long-range pairing interactions, where fermionic correlators decay algebraically even at non-critical points, signature of quantum criticality persists in the stationary state. Furthermore, the effective central charge, calculated from both mutual information and logarithmic negativity of stationary state following a critical-to-critical quench, agrees with the central charge of the corresponding ground states for both $\alpha=0$ and $\alpha=2$. Therefore, information of the universality class can be inferred from the stationary state.
I Introduction
Understanding quantum phase transitions (QPTs) has been a central topic of interest over the past few decades due to its significant implications for understanding the collective behavior of many-body systems [1]. QPTs have been both theoretically and experimentally studied extensively in various systems, including quantum spin chains [2; 3; 4; 5; 6; 7; 1; 8; 9; 10; 11; 12], Bose-Einstein condensates [13; 14; 15; 16; 17; 18; 19; 20], and strongly correlated electron systems [21; 22; 23; 24; 25]. QPT occurs at specific values of the parameter(s) present in Hamiltonian, called critical points where the system exhibits scale invariance. A universal behavior ensues due to the scale invariance, which can be classified into certain universality classes based on the space dimension and symmetry of the order parameter[26; 27; 28; 29]. This universality is studied using conformal field theories, and the corresponding central charge value is commonly used to identify the underlying universality class [30; 31; 32]. For instance, systems belonging to the Ising universality class are described by a minimal model with a central charge $c=1/2$, while the Luttinger liquid universality class corresponds to a conformal bosonic theory with $c=1$ [33; 34]. Note that at quantum critical points, the development of long-range correlations leads to an algebraic decay of correlators over distance, in contrast to the exponential decay typically observed at noncritical points in Hamiltonians with only short-range interactions [35; 36; 37; 38]. This is often used to identify critical points.
While the existence of quantum criticality and long-range correlations in the ground states of many-body quantum systems are widely studied, their presence in stationary states are not yet well explored. In Ref. [39], it is shown that when two sides of a 1D non-interacting fermionic chain with nearest-neighbor hopping are prepared at different temperatures, the mutual information scales logarithmically with subsystem size in the steady state, indicating the presence of long-range correlations. Another example of the survival of quantum criticality in the long-time stationary state is shown in the anisotropic XY chain under a sudden quench protocol [40]. Using quantum information measures such as mutual information and logarithmic(log-) negativity, it is shown that when both the pre- and post-quench parameters of the Hamiltonian are set at the critical point, both mutual information and log-negativity exhibit a peak, indicating the presence of criticality in the stationary state. In contrast, the peak in mutual information and log-negativity vanishes for other quench protocols when either/both pre- and post-quench parameters differ from critical values. The signature of quantum criticality in the stationary state is attributed to a change in correlation pattern from exponential to algebraic decay, which was captured by mutual information and log-negativity.
A natural question follows about the scenario when correlators decay algebraically, even at noncritical points? Long-range interacting systems provide an ideal platform to explore this question as fermionic correlators have been shown to decay algebraically over distance, even when the Hamiltonian is gapped [41; 42; 43; 44], i.e. at noncritical points. These long-range interacting systems have been studied in different contexts both theoretically and experimentally in atomic, molecular, and optical lattice systems [45; 46; 47; 48; 49; 50; 51; 52; 53; 54; 55; 56; 57; 58; 59; 60; 61]. We take the 1D long-range Kitaev model with a pairing term that decays with distance as a power law $\propto l^{-\alpha}$ with $\alpha$ being the exponent[44]. From an experimental perspective, this model is particularly relevant as it is closely related to the Ising model with tunable long-range interactions, which can be experimentally realized using trapped ion setups [53; 54; 55; 62; 58; 57]. In the ground state, this model undergoes an exotic transition from the Ising type universality class, observed for $\alpha>3/2$, to a Luttinger liquid universality class at $\alpha=0$.
In this paper, we investigate whether it is possible to capture quantum criticality in the stationary state when fermionic correlators decay algebraically even at non-critical points. Additionally, we examine whether the stationary state can be described by the same universality class as the ground state by analyzing the central charge obtained via scaling of mutual information and log-negativity for various quench protocols. Specifically, we consider three values of $\alpha$ corresponding to different universality classes observed in the ground state: first, corresponding to the Ising universality class ($\alpha=2$); second, the Luttinger liquid universality class ($\alpha=0$); and a third value ($\alpha=1$) where there is no universality [44].
II The model and ground state phase transition
The Hamiltonian of the $1$d long-range Kitaev model (LRK) for a lattice site of length $N$ is expressed as [44]
$H_{\rm LRK}=$ $\sum_{j=1}^{N}\left[-t\left(f_{j}^{\dagger}f_{j+1}+f_{j+1}^{\dagger}f_{j}\right)-\mu\left(f_{j}^{\dagger}f_{j}-\frac{1}{2}\right)\right.$ (1)
$\left.+\frac{\Delta}{2}\sum_{l=1}^{N-1}\frac{1}{l^{\alpha}}\left(f_{j+l}^{\dagger}f_{j}^{\dagger}-f_{j+l}f_{j}\right)\right]\,$
where $f_{j}(f_{j}^{\dagger})$ is the fermionic annihilation (creation) operator at site $j$, satisfying the canonical anticommutation relations $\{f_{i},f_{j}^{\dagger}\}=\delta_{ij}$, $\{f_{i},f_{j}\}=\{f_{i}^{\dagger},f_{j}^{\dagger}\}=0$. The parameters $t$ and $\mu$ represent the tunneling rate between two neighboring sites and the chemical potential, respectively. We set $12t=$ 1 in the rest of the paper. The parameter, $\Delta$ denotes the strength of the fermion $p$-wave pairing interaction, while its range is governed by the exponent $\alpha\in[0,\infty)$. The two limits, i.e. $\alpha=0$ and $\alpha\to\infty$ correspond to all-to-all interaction with equal strength and nearest-neighbour interaction, respectively. We consider the anti-periodic boundary condition throughout the paper. The Hamiltonian in Eq. 1 is exactly solvable. To see this, we first perform a Fourier transformation to transform Eq. 1 in the momentum space as
$H_{\rm LRK}=\frac{1}{2}$ $\sum_{n=0}^{N-1}\begin{bmatrix}f_{k_{n}}^{\dagger}&f_{N-k_{n}}\end{bmatrix}$ (2)
$\begin{bmatrix}-\left(\mu+\cos k_{n}\right)&i\Delta g_{\alpha}(k_{n})\\ -i\Delta g_{\alpha}(k_{n})&\left(\mu+\cos k_{n}\right)\end{bmatrix}\begin{bmatrix}f_{k_{n}}\\ f_{N-k_{n}}^{\dagger}\end{bmatrix}\,,$
where $g_{\alpha}(k)=\sum_{l=1}^{N-1}\frac{\sin(kl)}{l^{\alpha}}$ and $k_{n}=\frac{2\pi}{N}\left(n+1/2\right)$. In the large $N$ limit, $g_{\alpha}(k)$ takes the form
$g_{\alpha}(k)=-\frac{i}{2}\left[\text{Li}_{\alpha}\left(e^{ik}\right)-\text{Li}_{\alpha}\left(e^{-ik}\right)\right]\,.$ (3)
The function $\rm Li_{\alpha}(z)$ represents the polylogarithm of the complex variable $z$ of order $\alpha$. The Hamiltonian in Eq. 2 can further be cast to a diagonal form by performing the following Bogoliubov transformation
$\begin{bmatrix}f_{k_{n}}\\ f_{N-k_{n}}^{\dagger}\end{bmatrix}=\begin{bmatrix}\cos\theta_{k_{n}}&-i\sin\theta_{k_{n}}\\ -i\sin\theta_{k_{n}}&\cos\theta_{k_{n}}\end{bmatrix}\begin{bmatrix}\eta_{k_{n}}\\ \eta_{N-k_{n}}^{\dagger}\end{bmatrix},$ (4)
with $\theta$ being the Bogoliubov angle and is defined as
$\tan(2\theta_{k_{n}})=\frac{\Delta g_{\alpha}(k_{n})}{\mu+\cos k_{n}}.$ (5)
The diagonal form of the Hamiltonian in Eq. 1 in the basis of Bogoliubov fermions $\eta$ then becomes
$H_{\rm LRK}=\sum_{n=0}^{N-1}\lambda_{\alpha}(k_{n})\left(\eta_{k_{n}}^{\dagger}\eta_{k_{n}}-\frac{1}{2}\right),$ (6)
where the dispersion relation is
$\lambda_{\alpha}(k_{n})=\sqrt{\left(\mu+\cos k_{n}\right)^{2}+\left(\Delta g_{\alpha}(k_{n})\right)^{2}}.$ (7)
Equation 6 signifies that each modes are independent implying the integrability of the system. In the limit $\alpha\to\infty$, only nearest neighbour terms contribute to the sum in $g_{\alpha}(k)$, resulting in $g_{\infty}(k)=\sin(k)$. The Hamiltonian in Eq. 1 in this limit coincides with the Hamiltonian of the $XY$ model obtained via the Jordan-Wigner transformation. For the $XY$ model, a ground state phase transition occurs from the gapped ferromagnetic ordered phase to the gapped paramagnetic disordered phase with increasing the chemical potential $|\mu|$ with the two critical gapless points at $\mu=\pm 1$ [44]. The spectrum in Eq. 7 continues to be gapless at $\mu=\pm 1$ even with the smaller values of $\alpha$, i.e., in the presence of long-range interactions as long as $\alpha>1$. However, the phase diagram changes for $\alpha<1$. This regime is referred to as “strong” long-range regime since $\alpha<d$, with $d$ representing the dimension of the system. In comparison to $\alpha>1$, one significant change is that $\mu=-1$ is no longer a critical point, and the phase diagram is no longer symmetric across the line $\mu=0$ [44; 42].
Table 1: Summary of the effective central charge $c_{\rm eff}$ for different values of $\alpha$ and $\mu$ for ground state of long-range Kitaev Model [44; 42; 63; 64].
$\alpha=0$ $0<\alpha\leq 1$ $\alpha\to\infty$
$\mu=1$ 1 $c_{\text{eff}}(\alpha,\Delta)\neq 0$ $\dfrac{1}{2}$
$\mu\neq\pm 1$ $\dfrac{1}{2}$ $c_{\text{eff}}(\alpha,\Delta)\neq 0$ 0
A good probe to identify these critical points is the von Neumann entropy, $S_{vN}$, expressed as $S_{vN}=-{\text{Tr}}\left(\rho_{L}\ln\rho_{L}\right)$, where $\rho_{L}$ is the reduced density matrix of the subsystem of length $L$. At these critical points, where the spectral gap closes, $S_{vN}$ for short-range interacting system scales as [65; 66; 67; 68],
$S_{vN}\approx\frac{c}{3}\ln\left[\frac{N}{\pi}\sin\left(\frac{\pi L}{N}\right)\right]+c^{\prime},$ (8)
where $c$ is the central charge of the underlying conformal field theory (CFT), $N$ is the total system size and $c^{\prime}$ is a non-universal constant. In fact, Eq. 8 is also valid for all values of $\alpha$ and $\mu$, but the central charge $c$ is now replaced with an effective central charge $c_{\rm eff}$ [44; 63]. The values of the $c_{\rm eff}$ for ground state are summarized in Table-1. It is worth mentioning that $\alpha=0$ signifies all-to-all coupling between the sites and falls under the Tomonaga-Luttinger liquid universality class [69].
III Methods
In last section, we summarized criticality seen in ground state of LRK model and (effective) central charge of the underlying CFT. A natural question then arises whether this criticality survives when system undergoes out of equilibrium dynamics? In other words, in a stationary state can we find any signature of quantum criticality? A common way to achieve the non-equilibrium dynamics, is through sudden quench where the local or global parameters of the system are suddenly changed [70]. As a consequence, the ground state of the pre-quench Hamiltonian is no longer a ground state of the post-quench Hamiltonian but a superposition of it’s eigenstates. The long-time stationary state of the LRK model under sudden quench is the system of interest. To probe the criticality, much like ground state, entanglement entropy would be our first choice. However, as the long-time stationary state typically involves highly excited states, $S_{vN}$ exhibits volume-law scaling eclipsing the logarithmic dependence on subsystem size which may be present due to criticality. So, to extract these logarithmic correlations, we consider quantum information theoretic measures such as mutual information and log-negativity. For ground state, mutual information has already been established to capture the criticality in fermionic systems [71].
Mutual information: The mutual information between two subsystems $A_{1}$ and $A_{2}$ is defined as
$I_{A_{1}:A_{2}}=S_{vN}^{A_{1}}+S_{vN}^{A_{2}}-S_{vN}^{A_{1}\cup A_{2}},$ (9)
where $S_{vN}^{A_{1}}(S_{vN}^{A_{2}})$ is the von Neumann entropy of the subsystem $A_{1}(A_{2})$ and $S_{vN}^{A_{1}\cup A_{2}}$ is the von Neumann entropy of $A_{1}\cup A_{2}$. Throughout the paper, we consider the subsystem sizes $|A_{1}|=|A_{2}|=L$. The mutual information between two subsystems measures the total amount of information (both classical and quantum) that one subsystem contains about the other [72]. Furthermore, if $S_{vN}$ of the long time stationary state contains both the volume and logarithmic term, the mutual information should scale as $I_{A_{1}:A_{2}}\sim b\ln L$. Thus, if the long-range correlations are present in the stationary state, then the mutual information should behave as [73; 74; 75; 76; 39; 40]
$I_{A_{1}:A_{2}}\sim\frac{c_{\rm eff}^{I}}{3}\ln L+\rm const,$ (10)
where we define $c_{\rm eff}^{I}$ as the effective central charge of the stationary state extracted using mutual information. In this way, mutual information can capture the logarithmic correlation that is hidden in $S_{vN}$. However, whether these correlations are truly quantum in nature remains to be investigated. In this context, log-negativity serves as a useful measure.
Log-negativity: Similar to mutual information, log-negativity, an entanglement monotone, also quantifies the non-local correlation between two subsystems. But unlike mutual information, it captures only quantum correlations [77; 78; 79; 80]. It is defined as [78; 77]
$\xi_{A_{1}:A_{2}}\equiv\ln||\rho_{A}^{T_{2}}||=\ln{\text{Tr}}|\rho_{A}^{T_{2}}|,$ (11)
where ||.||||.||| | . | | denotes the trace norm, and $\rho_{A}^{T_{2}}$ signifies the partial transposition of the reduced density matrix of $A\equiv(A_{1}\cup A_{2})$. It is worth mentioning that for a CFT, the expression of log-negativity for two adjacent subsystems each of length $L_{1}$ and $L_{2}$, respectively, is shown to be [81]
$\xi_{A_{1}:A_{2}}=\frac{c}{4}\ln\left(\frac{L_{1}L_{2}}{L_{1}+L_{2}}\right),$ (12)
where $c$ is the central charge. For the $XY$ model, during a critical-to-critical quench, the log negativity of the long-time stationary state scales as $\xi_{A_{1}:A_{2}}=\frac{1}{8}\ln L$, where $L_{1}=L_{2}=L$ [40]. The logarithmic divergence of $\xi_{A_{1}:A_{2}}$ is consistent with the critical ground state, as the $XY$ model belongs to the Ising universality class with $c=1/2$. Based on this, we assume that if there is any signature of long-range quantum correlations in the long time stationary state, the log negativity would exhibit the following scaling behaviour:
$\xi_{A_{1}:A_{2}}\sim\frac{c_{\rm eff}^{N}}{4}\ln L+\rm const,$ (13)
where $c_{\rm eff}^{N}$ represents the effective central charge of the stationary state extracted using the log negativity.
Time-evolution: For non-interacting free-fermionic chains, owing to the simple relationship between the eigenvalues of reduced density matrix and two-point correlation function, we can calculate the entanglement entropy more efficiently by defining correlation matrix $W$ as,
$\begin{split}W_{nm}&={\text{Tr}}\left(\rho\begin{bmatrix}f_{n}\\ f_{n}^{\dagger}\end{bmatrix}\begin{bmatrix}f_{m}^{\dagger}&f_{m}\end{bmatrix}\right)\\ &=\begin{bmatrix}\delta_{nm}-C_{nm}&F_{nm}\\ \overline{F_{nm}}&C_{nm}\end{bmatrix},\end{split}$ (14)
[…]
$S=-\frac{1}{2}{\text{Tr}}\left[(I-W)\log(I-W)+W\log W\right].$ (15)
At any finite time $t$, the correlation functions are given by $C_{nm}(t)=\langle f_{n}^{\dagger}(t)f_{m}(t)\rangle$ and $F_{nm}(t)=\langle f_{n}(t)f_{m}(t)\rangle$, where the expectation values are taken with respect to the time evolved state obtained under the operation of the post-quench Hamiltonian. The two time-dependent correlation functions, $C_{nm}(t)$ and $F_{nm}(t)$ are expressed as [83]
$\langle f_{j}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{ik_{n}l}\bigg{[}-\sin\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\sin\left(2\delta\theta_{k_{n}}\right)$ (16)
$\quad+i\bigg{(}\sin\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\qquad-\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\theta_{k_{n}}\right)\bigg{)}\bigg{]}$
$\langle f_{j}^{\dagger}(t)f_{j+l}(t)\rangle$ $=\frac{1}{2N}\sum_{n=0}^{N-1}e^{-ik_{n}l}\bigg{[}1-\cos\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)$
$\quad-\sin\left(2\theta_{k_{n}}\right)\sin\left(2\delta\theta_{k_{n}}\right)\cos\left(2\lambda^{f}_{\alpha}(k_{n})t\right)\bigg{]},$
where $\delta\theta_{k_{n}}=\theta^{f}_{k_{n}}-\theta^{i}_{k_{n}}$ is the difference between pre- and post-quench Bogoliubov angles defined in Eq. 5 and $\lambda^{f}_{\alpha}(k_{n})$ is defined by Eq. 7 for post-quench Hamiltonian.
In the limit $t\to\infty$, the time dependent sine and cosine functions become highly oscillatory and the respective sum would tend to zero. Using this approximation, the expression for the correlation functions in the long-time stationary state is given by,
$\langle f_{j}f_{j+l}\rangle_{\text{st}}&=\frac{i}{2N}\sum_{n=0}^{N-1}e^{ik_{n}l}\sin\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)\\ \langle f_{j}^{\dagger}f_{j+l}\rangle_{\text{st}}&=\frac{1}{2N}\sum_{n=0}^{N-1}e^{ik_{n}l}\left[1-\cos\left(2\theta_{k_{n}}\right)\cos\left(2\delta\theta_{k_{n}}\right)\right],\end{aligned}$ (17)
where $\langle f_{j}^{\dagger}f_{j+l}\rangle_{\text{st}}$ and $\langle f_{j}f_{j+l}\rangle_{\text{st}}$ are two point correlation and two-point anomalous correlation function for stationary state. We utilize the above two expressions to generate the correlation matrix $W$, as defined by Eq. 14, and then apply Eq. 15 and Eq. 9 to obtain $S_{vN}$ and $I_{A_{1}:A_{2}}$, respectively, for the long-time stationary state.
Numerical estimation of log negativity: Equation 11 suggests that to evaluate log negativity, the crucial part is to perform the partial transposition. Unfortunately, the partial transposition of a fermionic Gaussian state is not a Gaussian state [84; 85]. This makes it difficult to efficiently calculate log-negativity in non-interacting fermionic systems. However, by expressing the partial transposition as a linear combination of two Gaussian operators, an upper bound on log-negativity can be obtained and is expressed as [85; 84; 86; 87; 88; 40; 89]
$\xi^{u}_{A_{1}:A_{2}}=\ln{\text{Tr}}\left(O_{+}O_{-}\right)^{\frac{1}{2}}+\ln\sqrt{2},$ (18)
where trace norm of $O_{+}$ is given by
$||O_{+}||$ $={\text{Tr}}\left(O_{+}O_{-}\right)^{\frac{1}{2}}$ (19)
$=\det\left[\left(\frac{\mathbb{I}+i\Gamma_{x}}{2}\right)^{\frac{1}{2}}+\left(\frac{\mathbb{I}-i\Gamma_{x}}{2}\right)^{\frac{1}{2}}\right]$
×$\quad\times\det\left(\frac{\mathbb{I}-\tilde{\Gamma}_{1}\tilde{\Gamma}_{2}}{2}\right),$
with
$\Gamma_{x}=i\left[1-\left(1+i\tilde{\Gamma}_{2}\right)\left(1-\tilde{\Gamma}_{1}\tilde{\Gamma}_{2}\right)^{-1}\left(1+i\tilde{\Gamma}_{1}\right)\right]\,.$ (20)
The quantities $\tilde{\Gamma}_{1},\tilde{\Gamma}_{2}$ in Eqs. 19 and 20 are defined as
$\tilde{\Gamma}_{k}=\tilde{M}_{2}\Gamma_{k}\tilde{M}_{2}\,,\quad\quad M_{2}=\begin{bmatrix}1\!\!1_{L}&0\\ 0&i1\!\!1_{L}\end{bmatrix}\,,$ (21)
and the size of the subsystems $|A_{1}|=|A_{2}|=L$. The correlation matrices $\Gamma_{1}$ and $\Gamma_{2}$ are defined as
$\Gamma_{1}=\Gamma^{A}\quad\Gamma_{2}=M_{2}\Gamma_{1}M_{2}\quad M_{2}=\begin{bmatrix}1\!\!1_{L}&0\\ 0&-1\!\!1_{L}\end{bmatrix},$ (22)
where $\Gamma^{A}$ is the correlation matrix corresponding to the subsystem $A$ ($A=A_{1}\cup A_{2}$) and is defined as $\Gamma^{A}=2W-I$. With Eqs. 19, 21 and 22, the upper bound of log negativity can be evaluated using Eq. 18.
Figure 1: In the first row, we plot the mutual information between two subsystems, $I_{A_{1}:A_{2}}$, of the stationary state as a function of the post-quench chemical potential $\mu_{f}=\mu$, for three values of $\alpha$: $\alpha=0$ (black squares), $\alpha=1$ (magenta stars), and $\alpha=2$ (orange pluses). In the left column, the pre-quench chemical potential $\mu_{i}$ is set at the critical point, i.e., $\mu_{i}=1$, while in the right column it is set at a non-critical point, with $\mu_{i}=1.5$. In the second row, we plot the log-negativity of the stationary state using the same symbols and quench parameters as in the first row. In the third row, we plot the occupation probability $n_{k}$ of Bogoliubov fermions at the soft mode, calculated from Eq. 23, with the same symbols and quench parameters as in the first row. For all figures, the quench protocol for the strength of the $p$-wave pairing interaction $\Delta$ is chosen as $\Delta_{i}=-1$ and $\Delta_{f}=1$.
IV Results
IV.1 Signature of criticality in stationary state
As we look for the signature of quantum criticality in the long-time stationary state for different $\alpha$ values, we consider two types of quench protocols which differ in the initial states. In the first protocol, we consider initial state as the ground state of LRK Hamiltonian corresponding to the parameter values $\mu_{i}=1$ and $\Delta_{i}=-1$ and will be referred henceforth as critical state. The second protocol correspond to the initial state as ground state of LRK Hamiltonian corresponding to parameters $\mu_{i}=1.5$ and $\Delta_{i}=-1$ which will be referred henceforth as non-critical state. The final Hamiltonian parameters, $\mu_{f}$ and $\Delta_{f}$, are varied to include both the critical and non-critical regimes. Note that it has already been established that mutual information of ground state shows a peak at the critical point namely $\mu=\pm 1$ for large $\alpha(=10)$ while only at $\mu=1$ for $\alpha\leq 1$ [71]. Mutual information, $I_{A_{1}:A_{2}}$ calculated for stationary state for both quench protocols i.e. $(\mu_{i}=1,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$ and $(\mu_{i}=1.5,\Delta_{i}=-1)\to(\mu_{f}=\mu,\Delta_{f}=1)$ are plotted in Fig.1 (row 1). For final quench parameters corresponding to critical values, a peak in $I_{A_{1}:A_{2}}$ for critical initial state is clearly visible indicating the signature of the criticality in stationary state for all values of $\alpha$. In the other quench protocol with non-critical initial state, $I_{A_{1}:A_{2}}$ does show a non-analytic behaviour in the form of a sharp dip for $\mu_{f}=1$ for all values of $\alpha$. To confirm that this behaviour is indeed due to quantum nature of correlations, we plot log-negativity $\xi_{A_{1}:A_{2}}$ in Eq. 11 in Fig. 1 (row 2) for the same quench protocols. The similar peak and dip structure is unmistakably seen in behaviour of $\xi_{A_{1}:A_{2}}$ whenever the final Hamiltonian is critical. Such behaviour of $I_{A_{1}:A_{2}}$ for nearest neighbour free-fermionic model has earlier been explained in terms of soft mode for which the energy vanishes near criticality. For a critical-to-critical quench, the soft mode does not get excited resulting in a peak structure in $I_{A_{1}:A_{2}}$ while it heats up to infinite temperature for a noncritical-to-critical quench displaying a dip in $I_{A_{1}:A_{2}}$. This further manifests as an algebraic decay of fermionic correlation for critical-to-critical quench, akin to ground-state behavior, while for the non-critical-to-critical quench, it decays exponentially indicating a non-critical behavior. However, such an explanation falls short for LRK model for $\alpha\leq 1$, as correlation decay pattern in the ground state is always algebraic.
For this non-analytic behaviour of entanglement measure at $\mu_{f}=1$, we offer an alternate explanation in terms of the mode occupation probability, $n_{k}$, for soft mode ($k_{c}=\pi$) where the energy vanishes near criticality. Let’s recall that $n_{k_{c}}=1/2$ would imply contribution from multiple modes and therefore leads to higher entanglement entropy for union of two subsystems, $S_{vN}^{A_{1}\cup A_{2}}$. As in $I_{A_{1}:A_{2}}$ in Eq. 9, $S_{vN}^{A_{1}\cup A_{2}}$ is being subtracted, this leads to decrease in $I_{A_{1}:A_{2}}$. In contrast, $n_{k_{c}}=1,0$ will lead to reduction of $S_{vN}^{A_{1}\cup A_{2}}$ and consequently leads to increment of the mutual information.
The occupation probability, $n_{k}=\langle\eta_{k}^{\dagger}\eta_{k}\rangle$ of the Bogoliubov modes, $\eta_{k}$, in the stationary state of LRK model is obtained as
$\langle\eta_{k}^{\dagger}\eta_{k}\rangle=\frac{1}{2}\left[1-\cos\left(2\delta\theta_{k}\right)\right]$ (23)
$=\frac{1}{2}\left[1-\frac{\left(\mu_{f}+\cos k\right)\left(\mu_{i}+\cos k\right)+\Delta_{f}\Delta_{i}g^{2}_{\alpha}\left(k\right)}{\lambda^{f}_{\alpha}\left(k\right)\lambda^{i}_{\alpha}\left(k\right)}\right],$
where $\lambda^{i}_{\alpha}\left(k\right)(\lambda^{f}_{\alpha}\left(k\right))$ denotes the energy corresponding to the pre(post)-quench Hamiltonian. From Eq. 23, it can be shown that if we start with the non-critical initial state, i.e., $\mu_{i}=1.5$, in the limit $k\to\pi$, the occupation probability. $n_{k}$ is given by
$n_{k}=\begin{cases}1&\text{for }\mu<1,\\ \frac{1}{2}&\text{for }\mu=1,\\ 0&\text{for }\mu>1,\end{cases}$ (24)
as illustrated in Fig.1(row 3). This explains the dips observed in $I_{A_{1},A_{2}}$ in the long-time stationary state for a non-critical-to-critical quench. In contrast, for the quench protocol involving critical state, $n_{k}$ near the soft mode is
$n_{k}=\begin{cases}\frac{1}{2}&\text{for }\mu<1,\\ 1&\text{for }\mu=1,\\ \frac{1}{2}&\text{for }\mu>1.\end{cases}$ (25)
This explains the peak observed in $I_{A_{1},A_{2}}$ for the critical-to-critical quench protocol.
Figure 2: Phase plot of effective central charge $c_{\rm eff}^{I}$ extracted from mutual information in the post-quench $\Delta-\mu$ plane for both critical (first column) and non-critical (second column) initial state for $\alpha=0$ (first row), 1 (second row) and 2 (third row). The parameter $\Delta_{i}$ is fixed at -1 for all the plots.
IV.2 Signature of universality in stationary state
After establishing the non-analytic behaviour of mutual information (and log-negativity) for post-quench stationary state at critical point of post-quench Hamiltonian, a natural question arises whether information of effective central charge, $c_{\rm eff}^{I}$ ($c_{\rm eff}^{N}$) calculated from scaling of $I_{A_{1}:A_{2}}$ ($\xi_{A_{1}:A_{2}}$) depends on the quench protocol? Whether this effective central charge is universal i.e.independent of quench protocols in $\Delta$, if yes whether the central charge of ground state is conserved? To address these questions, we analyze both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ for both the quench protocols discussed in sub-section IV.1. Here, it is useful to recall that central charge provides a powerful tool to understand the underlying CFT that describes the particular universality class [33]. For instance, the Ising universality class and the Luttinger liquid universality class are described by completely different CFT and characterized by distinct central charges. The central charge gives a measure of the number of degrees of freedom or the “size” of the symmetry in the theory [90; 91].
Fig. 2 displays $c^{I}_{\rm eff}$ as a function of the post-quench parameters $\mu_{f}$ and $\Delta_{f}$. The columns of Fig. 2 represents $c^{I}_{\rm eff}$ for quenches starting with two distinct initial states, while rows correspond to varying range of $\alpha$. Let’s focus on the first column i.e. quench from critical state for different values of $\alpha$. For both $\alpha=0$ and $\alpha=2$, $c_{\rm eff}^{I}$ is independent of $\Delta$-quenches and retain the same value as the ground state for a critical-to-critical quench protocol. For other values of $\mu_{f}$, $c_{\rm eff}^{I}$ calculated from stationary state takes the value of non-critical ground state, i.e. $0.5$ for $\alpha=0$ and 0 for $\alpha=2$  [^1] Note that $\Delta=0$ completely changes the Hamiltonian to a nearest neighbour model with no pairing term with effective central charge 0 for $\mu\neq 1$.. For $\alpha=1$, the effective central charge for ground state at critical point depends on the value of $\Delta$ and this character is retained in the post-quench stationary state as well. Therefore, the universality can only be discussed in the case of $\alpha=0$ and $\alpha=2$ from a critical-to-critical quench. It is clearly seen that $c_{\rm eff}^{I}$ calculated from post-quench stationary state conserves the central charge (see Table-1). This implies that universality class for the critical ground state can also be inferred from stationary state.
Table 2: Summary of effective central charge values calculated from post-quench stationary state for critical-to-critical, noncritical-to-critical and critical-to-noncritical quenches of LRK model for different values of $\alpha$.
$\alpha=0$ $0<\alpha\leq 1$ $\alpha=2$
$\mu_{i}=1,\mu_{f}=1$ 1 $c_{\text{eff}}(\alpha,\Delta)\neq 0$ $\dfrac{1}{2}$
$\mu_{i}\neq 1,\mu_{f}=1$ $\dfrac{1}{2}$ $c_{\text{eff}}(\alpha,\Delta)\neq 0$ 0
$\mu_{i}=1,\mu_{f}\neq 1$ $\dfrac{1}{2}$ $c_{\text{eff}}(\alpha,\Delta)\neq 0$ 0
Now for the second column of Fig. 2, universal characteristic of the $c_{\rm eff}^{I}$ of stationary state is reflected through independence of $\Delta$-quenches for $\alpha=0$ and $\alpha=2$ while $\alpha=1$ remains the $\Delta$-dependent. Note that even though entanglement measures showed a non-analytic behaviour at $\mu_{f}=1$ for non-critical initial state, the effective central charge does not show any such transition for either $\alpha=0$ or 2. Therefore, despite the non-analytic behaviour of entanglement measures at $\mu_{f}=1$, we do not attribute as signature of criticality or an universal description in terms of underlying CFT.
Therefore, to summarize, signature of universality in terms of effective central charge is seen in post-quench stationary state for a critical-to-critical quench. The effective central charges values in this case is summarized in tabular form in Table 2 which are in agreement with the values in Table 1 calculated for ground state.
Now, to investigate whether the long-range correlations in the stationary state are genuinely quantum in nature, we have evaluated the effective central charge, $c_{\rm eff}^{N}$, extracted from the finite-size scaling of the log-negativity for the stationary state. While the details are in Appendix A, the findings of effective central charge behaviour and therefore universality remains same as calculated using mutual information.
V Summary
In summary, we studied the quench dynamics of the long-range Kitaev chain to investigate the possible signatures of quantum criticality in the long-time stationary state for different values of the exponent $\alpha$ characterizing the range of $p$-wave pairing interaction. To probe the criticality, we consider the quantum information-based measures such as mutual information and log-negativity. Our results show that a peak emerges in both mutual information and log-negativity for a critical-to-critical quench for all values of $\alpha$ studied in this work, exhibiting behavior similar to that of the critical ground state. This clearly indicates the presence of criticality in the stationary state. Importantly, the appearance of this peak can not be attributed to the change in decay pattern of fermionic correlation from exponential to algebraic at the critical point, as the fermionic correlators decay algebraically even at non-critical points for $\alpha\leq 1$. We argue that the peak arises as a result of the change in the occupation probability of the Bogoliubov fermions in the soft mode at the critical point.
We further show that for $\alpha=2$, long-range correlations develop only for the critical-to-critical quench. This result is obtained by analyzing the effective central charges, which are extracted from the scaling analysis of both mutual information and log-negativity. The effective central charge of the corresponding stationary state matches with the same for the critical ground state, regardless of the quench in the strength of the $p$-wave pairing interaction, $\Delta$. This clearly indicates that universality class for the critical ground state can also be inferred from stationary state. In contrast, for $\alpha=0$, the effective central charge is non-zero for all quench protocols, leading to long-range correlations, except when the final Hamiltonian parameter, $\Delta=0$. When $\Delta\neq 0$, the effective central charge is $1$ only for the critical-to-critical quench, and is $1/2$ for all other quench protocols, consistent with the ground state. This infers that the stationary state for critical to critical quench protocol belongs to the same universality class as the ground state.
For $\alpha=1$, the stationary state, similar to the ground state, cannot be described by any universality class, as the presence or absence of long-range correlations depends on the $\Delta$ quenches. However, long-range correlations do develop for the critical-to-critical quench protocol, regardless of the $\Delta$ quenches.
VI Acknowledgment
S.P. would like to thank DST India for the Inspire Faculty Grant.
Appendix A Overview of central charge for post-quench stationary state from log-negativity
Starting with the critical initial state, as shown in first row left column of Fig. 3, we find that for $\alpha=0$, $c_{\rm eff}^{N}=1/2$ for $\mu_{f}\neq 1$ and $\Delta_{f}\neq 0$, while $c_{\rm eff}^{N}=1$ for $\mu_{f}=1$ and $\Delta_{f}\neq 0$. In contrast, starting with a non-critical initial state, as shown in first row right column of Fig. 3, $c_{\rm eff}^{N}=1/2$ for all values of $\mu_{f}$ and $\Delta_{f}$, except for $\Delta_{f}=0$. This implies that for $\alpha=0$, $c_{\rm eff}^{N}=c_{\rm eff}^{I}$ for all quench protocols. For $\alpha=2$, similar to $c_{\rm eff}^{I}$, $c_{\rm eff}^{N}$ is non-zero and equal to $1/2$ only for a critical-to-critical quench protocol (with the exception of $\Delta_{f}=0$, where $c_{\rm eff}^{N}=0$). For all other quench protocols, $c_{\rm eff}^{N}=0$. This confirms that the long-range correlations associated with a non-zero effective central charge for both $\alpha=0$ and $\alpha=2$ in the stationary state are strictly quantum in nature. On the other hand, for $\alpha=1$, long-range correlations develop irrespective of $\Delta$ quenches in the critical-to-critical quench protocol (see second row first column of Fig. 3). Interestingly, for this quench protocol, $c_{\rm eff}^{N}$ is slightly larger than $c_{\rm eff}^{I}$. In contrast, for the critical-to-non-critical quench protocol, the presence of long-range correlations in the stationary state depends on the $\Delta$ quenches. Specifically, if $\Delta_{f}\Delta_{i}\gtrsim 0$, both $c_{\rm eff}^{I}$ and $c_{\rm eff}^{N}$ are non-zero, whereas both are zero if $\Delta\Delta_{0}\lesssim 0$. For all the quench protocols where $\Delta_{f}\Delta_{i}\gtrsim 0$, $c_{\rm eff}^{N}>c_{\rm eff}^{I}$.
Figure 3: Phase plot of the effective central charge from log-negativity $c_{\rm eff}^{N}$ for both critical (first column) and non-critical (second column) initial state for $\alpha=0$ (first row), 1 (second row) and 2 (third row). The quench parameters are same as for Fig. 2.